Bryan Carpenter

313 East Genesee St, Apt 3 Fayetteville, NY 13066 USA

November 2, 1999

Macmillan Cancer Relief 15-19 Britten Street London, SW3 3BR

Dear Macmillan Cancer Relief,

Following the wishes of his wife Chris, this donation is in memory of my friend David Pritchard of 4, Westridge Road, Southampton, who died on October 25th, 1999. Please find a cheque enclosed.

Yours sincerely,